OpenStudy (jazzyluvsu123):

is 58.68/ 2 = 29.34 ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

58.68/ 2 = (50 + 8.68)/2 = 50/2 + 8.68/2 = 25 + 4.34 = 29.34

OpenStudy (jazzyluvsu123):

ok

OpenStudy (unklerhaukus):

58.68/ 2 = (50 + 8 + 0.6 + 0.08)/2 = 50/2 + 8/2 + 0.6/2 + 0.08 = 25 + 4 + 0.3 + 0.04 = 29.34